OpenStudy (clamin):

A data set consists of the following data points: (3,5) (5,8) (6,13) The slope of the best-fit line is 2.5. Find the y-intercept of this line. a. -3 b. -4 c. -5 d. -2

OpenStudy (anonymous):

The average x and y from (3,5) (5,8) (6,13) is\[\left\{\frac{14}{3},\frac{26}{3}\right\} \]\[y=b+2.5 x\text{ /.}\, \left\{x\to 4+\frac{2}{3},y\to 8+\frac{2}{3}\right\} //Solve\]b = -3.